OpenStudy (anonymous):

This problem addresses some common algebraic errors. For the equalities stated below assume that x and y stand for real numbers. Assume that any denominators are non-zero. Mark the equalities with T (true) if they are true for all values of x and y, and F (false) otherwise. (x+y)^2 = x^2 + y^2. (x+y)^2 = x^2 + 2xy+y^2. \frac{x}{ x+y} = \frac{1}{y}. x-(x+y) = y. \sqrt{x^2} = x. \sqrt{x^2} = |x|. \sqrt{x^2+4} = x+2. \frac{1}{x+y} = \frac{1}{x} + \frac{1}{y}.

OpenStudy (anonymous):

i got 1=F 2=T 3=F 4=T 5=T 6=T 7=F 8=T

OpenStudy (anonymous):

DONT NO WHICH ONES WRONG I SUBBED IN 2 FOR ALL THE X'S AND 4 FOR ALL THE Y'S

ganeshie8 (ganeshie8):

your answers for first 3 equalities are right

ganeshie8 (ganeshie8):

lets look at #4 : ` x-(x+y) = y`

ganeshie8 (ganeshie8):

distribute the negative for parenthesis on left hand side, what do u get ?

OpenStudy (anonymous):

do you no multiply brackets by -??

OpenStudy (anonymous):

i get -4 when i do that

ganeshie8 (ganeshie8):

yes you may think of : x-(x+y) as x - 1(x+y)

ganeshie8 (ganeshie8):

distributing gives you : x - 1x - 1y x - x - y 0 - y -y which is NOT same as the right hand side y eh ?

OpenStudy (anonymous):

woopsy your right that is false but i still i have another one or two of them wrong

ganeshie8 (ganeshie8):

yes lets see #5

ganeshie8 (ganeshie8):

\sqrt{x^2} = x

ganeshie8 (ganeshie8):

say x = -2 what do you get ?

ganeshie8 (ganeshie8):

sqrt((-2)^2) = -2 sqrt(4) = -2

ganeshie8 (ganeshie8):

can squareroot of something ever be negative ?

OpenStudy (anonymous):

no so therefore it is false?

ganeshie8 (ganeshie8):

it has to be false an equation has to work for all numbers

ganeshie8 (ganeshie8):

so what do we have so far ?

OpenStudy (anonymous):

o i see where your coming from now

OpenStudy (anonymous):

we have F T F F F

OpenStudy (anonymous):

the next 3 i have T F T

ganeshie8 (ganeshie8):

yes #6 is T

ganeshie8 (ganeshie8):

#7 is F

ganeshie8 (ganeshie8):

#8 is also F

ganeshie8 (ganeshie8):

so overall only #2 and #6 are True and everything else is False

OpenStudy (anonymous):

thank you very helpful

ganeshie8 (ganeshie8):

np :)